Marion needs to find the cosine of Pi over 12. If she knows that cos Pi over 6 = root 3 over 2, how can she use this fact to find the cosine of Pi over 12? What is her answer. Please explain and show your work
1
Expert's answer
2020-10-26T19:19:29-0400

Use double angle formula for cosine "cos(2A)=2cos^2(A)-1"


So, the cosine of angle "A" can be written as,


"2cos^2(A)=1+cos(2A)"


"cos^2(A)=\frac{1}{2}(1+cos(2A))"


"cos(A)=\sqrt{\frac{1}{2}(1+cos(2A))}"


Plug "A=\frac{\pi}{12}" into the relation "cos(A)=\sqrt{\frac{1}{2}(1+cos(2A))}" to obtain,


"cos(\frac{\pi}{12})=\sqrt{\frac{1}{2}(1+cos(2(\frac{\pi}{12})))}"


"=\sqrt{\frac{1}{2}(1+cos(\frac{\pi}{6}))}"


"=\sqrt{\frac{1}{2}(1+\frac{\sqrt{3}}{2})}" .....plug "cos(\frac{\pi}{6})=\frac{\sqrt3}{2}"


"=\sqrt{\frac{1}{2}(\frac{2+\sqrt{3}}{2})}"


"=\sqrt{\frac{1}{4}(2+\sqrt{3})}"


"=\frac{1}{2}\sqrt{2+\sqrt{3}}\approx0.965926"


Therefore, the cosine of "\frac{\pi}{12}" is "cos(\frac{\pi}{12})=\frac{1}{2}\sqrt{2+\sqrt{3}}\approx0.965926"